# Revisiting Node Affinity Prediction in Temporal Graphs

### v2

@@ Abstract
Node affinity prediction is a common task that is widely used in temporal graph learning with applications in social and financial networks, recommender systems, and more. Recent works have addressed this task by adapting state-of-the-art dynamic link property prediction models to node affinity prediction. However, simple heuristics, such as Persistent Forecast or Moving Average, outperform these models. In this work, we analyze the challenges in training current Temporal Graph Neural Networks for node affinity prediction and suggest appropriate solutions. Combining the solutions, we develop NAViS - Node Affinity prediction model using Virtual State, by exploiting the equivalence between heuristics and state space models. While promising, training NAViS is non-trivial. Therefore, we further introduce a novel loss function for node affinity prediction. We evaluate NAViS on TGB and show that it outperforms the state-of-the-art, including heuristics. Our source code is available at https://github.com/orfeld415/NAVIS
@@ 1 Introduction
Temporal graphs provide a natural way to represent evolving interactions in systems such as trade networks, recommender systems, social platforms, and financial transactions (Kumar et al., 2019; Shetty & Adibi, 2004; Huang et al., 2023). A central challenge in this setting is future node affinity prediction: forecasting how strongly a node will interact with other nodes at a future time. This differs from future link prediction, which instead asks whether a particular edge will appear. In contrast, affinity prediction requires producing a full ranking over potential neighbors, making it more demanding but also more relevant to many real-world applications (MacDonald et al., 2015; Bertin-Mahieux et al., 2011; Nadiri & Takes, 2022; Shamsi et al., 2022).
@@
Accordingly, this paper aims to answer the following questions: why do heuristics outperform more sophisticated TGNNs for future node affinity prediction? and can we push TGNNs to do better? We argue that the advantage arises from a confluence of factors and identify several contributing issues that collectively explain this phenomenon, including: (i) Expressivity.Existing TGNNs cannot represent a simple Moving Average of past affinities, because their nonlinear updates and reliance on sampled neighborhoods prevent them from maintaining the required linear memory. (ii) Loss mismatch.Cross-entropy, commonly used as a loss function for link prediction, is not well-aligned with the ranking nature of affinity tasks. (iii) Global temporal dynamics.Affinities often depend on shared network-wide trends (e.g., regime shifts), which local sampling does not capture. (iv) Information loss.TGNNs are broadly categorized into memory-based and non-memory-based architectures. Memory-based models (e.g., TGN (Rossi et al., 2020) and DyRep (Trivedi et al., 2019)) maintain per-node states; however, batch processing of events can cause short-term updates within a batch to be missed (Feldman & Baskin, 2024). In contrast, non-memory-based methods recompute embeddings from scratch at prediction time and maintain no states, making them prone to overlooking earlier updates in the evolving graph (Cong et al., 2023; Yu et al., 2023).
These observations guide and motivate our work. Our key idea is that heuristics like persistent Forecast and Moving Average are not arbitrary fixes, but rather special cases of linear state space models (SSMs) (Gu et al., 2021; 2022; Eliasof et al., 2025), which naturally provide memory and long-range temporal dependencies. By embedding the structure of SSMs into a learnable TGNN, we can retain the robustness of heuristics while extending their expressivity.
@@
To illustrate the existing challenges and the effectiveness of NAViS, we include a synthetic node-affinity experiment in Figure˜1. In this controlled setting, node affinities depend on a hidden global process with nonlinear structure. Although simple heuristics outperform state-of-the-art TGNNs, since they capture only per-node history, as TGNNs, they fail to recover the shared latent. By maintaining a virtual global state, NAViS achieves the lowest error. Although simplified, this example highlights the importance of global temporal dynamics and motivates one of the design choices in NAViS. We provide the full experiment details in Appendix˜C.
@@
Figure 1: Synthetic node affinity experiment. Node affinities depend on a global, regime-switching latent $g(t)$ with nonlinear component $g(t)^{2}$ and node-specific phases. Baselines relying only on per-node histories (Persistent Forecast, SMA, EMA) or a local AR(1) SSM cannot recover the shared latent, leading to higher error. Ours, indicating NAViS, maintains a virtual global state, achieves the lowest error on both metrics. In Appendix˜C we provide the full experiment details and baseline descriptions.
@@ 3.1 What Hinders TGNNs in Future Node Affinity Prediction
As established in Section˜2, common heuristics like PF and EMA are special cases of SSMs. This advantage of SSMs over heuristics is achieved due to the fact that SSMs can output a linear combination of the previous affinity vector and the previous predicted affinity vector. We now discuss another important theoretical direction: RNN (Elman, 1990), LSTM (Hochreiter & Schmidhuber, 1997) or GRU (Cho et al., 2014) cells — that are commonly used in popular memory-based TGNNs (Trivedi et al., 2019; Rossi et al., 2020; Tjandra et al., 2024) — cannot express the most basic heuristic of PF, thereby hindering memory-based TGNNs performance.
@@
Theorem˜2 has profound implications: any memory-based TGNN that applies standard memory cells cannot represent even the simplest heuristic – Persistent Forecasting (PF). While such models may be powerful for link prediction, they lack the functional expressivity required for node affinity prediction. This theoretical result is at the underpinnings of our work, and motivates us to generalize heuristics while remaining more expressive. Thus, we design NAViS as a simple learnable linear SSM.
@@
We argue that a major source of empirical underperformance in TGNNs is the systematic loss of temporal information. Memory–based TGNNs often rely on batching for tractable runtimes; however, batching can obscure multiple interactions that affect the same node within a single batch window, thereby dropping intermediate state transitions (Feldman & Baskin, 2024). In contrast, non–memory architectures, e.g., DyGFormer (Yu et al., 2023), GraphMixer (Cong et al., 2023), and DyGMamba (Ding et al., 2025), avoid within-batch omissions by maintaining a buffer of recent events and recomputing node embeddings on demand. To bound latency, these buffers are fixed in size; once filled, the oldest events are evicted, discarding potentially informative long-term interaction events. This hard truncation differs from EMA, where the influence of older events decays but remains non-zero, allowing previous interaction events to shape future affinity predictions. A further, underexploited source of information is the evolving global graph state. Memory–based TGNNs typically use few message-passing layers per state update, limiting the incorporation of broader context (Rossi et al., 2020), while non–memory methods commonly restrict buffered events to 1-hop neighborhoods. Even new sophisticated state-of-the-art TGNNs do not utilize the full global state (Lu et al., 2024; Gravina et al., 2024). We leverage these observations to design a TGNN that preserves fine-grained temporal transitions, retains long-term interaction events, and integrates global graph context to maximize the use of available information to give an accurate prediction for future node affinity.
@@ 3.2 NAViS: Node Affinity Prediction with a Global Virtual State
We compute the virtual global vector $\mathbf{g}$} by maintaining a buffer of the most recent previous affinity vectors, globally. Then, the virtual global vector is computed by performing aggregation over all the vectors in the buffer. In practice, aggregating the buffer with the most recent vector selection is efficient and empirically effective, as we show later in Section˜4.
@@
Key Properties of NAViS. We note that NAViS generalizes EMA and other heuristics by allowing the gates ${\mathbf{z}}_{h}$ and ${\mathbf{z}}_{s}$ to adapt to new information, in contrast to a fixed $\alpha$. Large gate values enable the model to retain long-term information when beneficial. Notably, NAViS does not rely on neighbors’ hidden states, unlike other memory-based TGNNs, and therefore is compatible with the t-Batch mechanism (Kumar et al., 2019), enabling efficient batching without missing updates. In addition, we show in Appendix˜E that incorporating global information via $\mathbf{g}$} can improve the accuracy of the predicted affinity vector when global trends affect the nodes’ affinities.
NAViS for large-scale graphs. For graphs with $N$ nodes, the number of learnable parameters in NAViS scales as $O(N)$, which can be prohibitive for graphs with millions of nodes. To make NAViS practical at this scale, we introduce a sparsified affinity prediction pipeline. Specifically, for each node we retain only the entries corresponding to candidate target nodes. In real-world settings, such as streaming services where users and movies are nodes and interactions are edges, we are interested in each user’s affinity to movies rather than to other users. In practice, this substantially reduces the parameter count of NAViS. For example, on the tgbn-token dataset (Shamsi et al., 2022), which records user–token interactions, NAViS requires about (5,000) parameters, while the graph contains over (60,000) nodes.
@@ 3.3 Learning with Rank-Based Loss: Why Cross-Entropy Fails
With NAViS specified in Section˜3.2, we now turn to the question of how to train it effectively. Because most downstream uses of affinity vectors depend on the induced ordering of candidates rather than the actual affinity values (Huang et al., 2023; Tjandra et al., 2024), the choice of the loss function is critical. Most TGNNs use the cross-entropy loss (Luo & Li, 2022; Yu et al., 2023; Tjandra et al., 2024), which treats the output as a categorical distribution and ignores ordinal structure. As we show next, this property is suboptimal, and requires a designated loss to address this limitation.
@@
There exist infinitely many triplets of $\mathbf{y}$}, a ground-truth affinity vector, and ${\mathbf{s}}_{1}$, ${\mathbf{s}}_{2}$, two predicted affinity vectors such that:
$\mathrm{rank}({\mathbf{s}}_{1})$ = $\mathrm{rank}({\mathbf{y}})$ and $\mathrm{rank}({\mathbf{s}}_{2})\neq\mathrm{rank}({\mathbf{y}})$, where $\ell_{\mathrm{CE}}({\mathbf{s}}_{1},{\mathbf{y}})>\ell_{\mathrm{CE}}({\mathbf{s}}_{2},{\mathbf{y}})$
@@
Set ${\mathbf{y}}=[0.4,0.35,0.25]$, ${\mathbf{s}}_{1}=[0.8,0.15,0.05]$ (correct rank), ${\mathbf{s}}_{2}=[0.35,0.4,0.25]$ (wrong rank). Then $1.091=\ell_{\mathrm{CE}}({\mathbf{s}}_{2},{\mathbf{y}})<\ell_{\mathrm{CE}}({\mathbf{s}}_{1},{\mathbf{y}})=1.105$, where ${\mathbf{s}}_{1}$ ranks the same as $\mathbf{y}$} while ${\mathbf{s}}_{2}$ does not. Since $\ell_{\mathrm{CE}}$ is a continuous function, there are infinitely many such triplets. ∎
@@
Pairwise Margin Regularization. In our experiments, we discovered that the use of the loss in Equation˜16 alone is not sufficient for future node affinity prediction tasks, as we elaborate in Appendix˜E. Hence, we suggest the following regularization:
@@
Here, $\Delta$ is a hyperparameter that represents the minimum margin required between each pair of affinity scores. The goal of this regularization is to prevent NAViS from incorrectly learning to shrink the affinity scores to minimize Equation˜16.
@@ 4 Experiments
We evaluate NAViS across multiple future node affinity prediction benchmarks and compare it with recent state-of-the-art baselines, including both heuristics and TGNNs. Sections˜4.1 and 4.2 present our key empirical findings, and Appendix˜E provides additional ablation studies. Our experiments aim to address the following research questions: (RQ1) How does NAViS perform compared to prior art for future node affinity prediction? (RQ2) Does our method generalize across various types of graphs? (RQ3) What is the contribution of each component in NAViS?
Experimental setup We compare NAViS to the following TGNN baselines JODIE(Kumar et al., 2019), TGAT(Xu et al., 2020), CAWN (Wang et al., 2021b), TCL (Wang et al., 2021a), GraphMixer(Cong et al., 2023),DyGFormer(Yu et al., 2023),DyRep(Trivedi et al., 2019), TGN(Rossi et al., 2020), TGNv2(Tjandra et al., 2024) and the standard heuristics presented in (Huang et al., 2023) and in (Tjandra et al., 2024). Following the standard protocols (Huang et al., 2023), we use a 70%-15%-15% chronological split, train for 50 epochs, use a batch size of 200, and report the average NDCG@10 (Järvelin & Kekäläinen, 2002) over three runs. We include both future node affinity prediction settings: (1) using the full fine-grained CTDG up to the prediction time, and (2) using only previous ground-truth affinity vectors. Previous TGNNs only support the first setting (Tjandra et al., 2024), and, therefore, only heuristics are included in the second setting comparisons.

### v3

@@ Abstract
Node affinity prediction is a common task that is widely used in temporal graph learning with applications in social and financial networks, recommender systems, and more. Recent works have addressed this task by adapting state-of-the-art dynamic link property prediction models to node affinity prediction. However, simple heuristics, such as Persistent Forecast or Moving Average, outperform these models. In this work, we analyze the challenges in training current Temporal Graph Neural Networks for node affinity prediction and suggest appropriate solutions. Combining the solutions, we develop NAViS - Node Affinity prediction model using Virtual State, by exploiting the equivalence between heuristics and state space models. While promising, training NAViS is non-trivial. Therefore, we further introduce a dedicated loss function for node affinity prediction. We evaluate NAViS on TGB and show that it outperforms the state-of-the-art, including heuristics. Our source code is available at https://github.com/orfeld415/NAVIS
@@ 1 Introduction
Temporal graphs provide a natural way to represent evolving interactions in systems such as trade networks, recommender systems, social platforms, and financial transactions (Kumar et al., 2019; Shetty and Adibi, 2004; Huang et al., 2023). A central challenge in this setting is future node affinity prediction: forecasting how strongly a node will interact with other nodes at a future time. This differs from future link prediction, which instead asks whether a particular edge will appear. In contrast, affinity prediction requires producing a full ranking over potential neighbors, making it more demanding but also more relevant to many real-world applications (MacDonald et al., 2015; Bertin-Mahieux et al., 2011; Nadiri and Takes, 2022; Shamsi et al., 2022).
@@
Accordingly, this paper aims to answer the following questions: why do heuristics outperform more sophisticated TGNNs for future node affinity prediction? and can we push TGNNs to do better? We argue that the advantage arises from a confluence of factors and identify several contributing issues that collectively explain this phenomenon, including: (i) Expressivity.Existing TGNNs cannot represent a simple Moving Average of past affinities, because their nonlinear updates and reliance on sampled neighborhoods prevent them from maintaining the required linear memory. (ii) Loss mismatch.Cross-entropy, commonly used as a loss function for link prediction, is not well-aligned with the ranking nature of affinity tasks. (iii) Global temporal dynamics.Affinities often depend on shared network-wide trends (e.g., regime shifts), which local sampling does not capture. (iv) Information loss.TGNNs are broadly categorized into memory-based and non-memory-based architectures. Memory-based models (e.g., TGN (Rossi et al., 2020) and DyRep (Trivedi et al., 2019)) maintain per-node states; however, batch processing of events can cause short-term updates within a batch to be missed (Feldman and Baskin, 2024). In contrast, non-memory-based methods recompute embeddings from scratch at prediction time and maintain no states, making them prone to overlooking earlier updates in the evolving graph (Cong et al., 2023; Yu et al., 2023).
These observations guide and motivate our work. Our key idea is that heuristics like persistent Forecast and Moving Average are not arbitrary fixes, but rather special cases of linear state space models (SSMs) (Gu et al., 2021; 2022; Ceni et al., 2025; Eliasof et al., 2025), which naturally provide memory and long-range temporal dependencies. By embedding the structure of SSMs into a learnable TGNN, we can retain the robustness of heuristics while extending their expressivity.
@@
To illustrate the existing challenges and the effectiveness of NAViS, we include a synthetic node-affinity experiment in Figure˜1. In this controlled setting, node affinities depend on a hidden global process with nonlinear structure. Although simple heuristics outperform state-of-the-art TGNNs, since they capture only per-node history, as TGNNs, they fail to recover the shared latent space. By maintaining a virtual global state, NAViS achieves the lowest error. Although simplified, this example highlights the importance of global temporal dynamics and motivates one of the design choices in NAViS. We provide the full experiment details in Appendix˜C.
@@
Figure 1: Synthetic node affinity experiment. Node affinities depend on a global, regime-switching latent $g(t)$ with nonlinear component $g(t)^{2}$ and node-specific phases. Baselines relying only on per-node histories (Persistent Forecast, SMA, EMA) or a local AR(1) SSM cannot recover the shared latent space, leading to higher error. Ours, indicating NAViS, maintains a virtual global state, achieves the lowest error on both metrics. In Appendix˜C we provide the full experiment details and baseline descriptions.
@@ 3.1 What Hinders TGNNs in Future Node Affinity Prediction
As established in Section˜2, common heuristics like PF and EMA are special cases of SSMs. This advantage of SSMs over heuristics is achieved due to the fact that SSMs can output a linear combination of the previous affinity vector and the previous predicted affinity vector. We now discuss another important theoretical direction: RNN (Elman, 1990), LSTM (Hochreiter and Schmidhuber, 1997) or GRU (Cho et al., 2014) cells — that are commonly used in popular memory-based TGNNs (Trivedi et al., 2019; Rossi et al., 2020; Tjandra et al., 2024) — cannot express the most basic heuristic of PF, thereby hindering memory-based TGNNs performance.
@@
Theorem˜2 has profound implications: any memory-based TGNN that applies standard memory cells cannot represent even the simplest heuristic – Persistent Forecasting (PF), which have been proven empirically to perform exceptionally well on node affinity tasks. This theoretical result is at the underpinnings of our work, and motivates us to generalize heuristics while remaining more expressive. Thus, we design NAViS as a simple learnable linear SSM. Tjandra et al. (2024) showed that identifying the target node when updating the node state of the source node is a necessary property. Without this property, TGNNs cannot express the persistent forecast heuristic. In the proof of Theorem˜2, we explicitly assume that the target node is identified via its corresponding index in the affinity vector of the source node. Hence, Theorem˜2 holds even when the target node is identified, revealing another necessary condition for expressing the persistent forecast heuristic.
@@
We argue that a major source of empirical underperformance in TGNNs is the systematic loss of temporal information. Memory–based TGNNs often rely on batching for tractable runtimes; however, batching can obscure multiple interactions that affect the same node within a single batch window, thereby dropping intermediate state transitions (Feldman and Baskin, 2024). In contrast, non–memory architectures, e.g., DyGFormer (Yu et al., 2023), GraphMixer (Cong et al., 2023), and DyGMamba (Ding et al., 2025), avoid within-batch omissions by maintaining a buffer of recent events and recomputing node embeddings on demand. To bound latency, these buffers are fixed in size; once filled, the oldest events are evicted, discarding potentially informative long-term interaction events. This hard truncation differs from EMA, where the influence of older events decays but remains non-zero, allowing previous interaction events to shape future affinity predictions. A further, underexploited source of information is the evolving global graph state. Memory–based TGNNs typically use few message-passing layers per state update, limiting the incorporation of broader context (Rossi et al., 2020), while non–memory methods commonly restrict buffered events to 1-hop neighborhoods. Notably, also recent sophisticated state-of-the-art TGNNs do not utilize the full global state (Lu et al., 2024; Gravina et al., 2024). We leverage these observations to design a TGNN that preserves fine-grained temporal transitions, retains long-term interaction events, and integrates global graph context to maximize the use of available information to give an accurate prediction for future node affinity.
@@ 3.2 NAViS: Node Affinity Prediction with a Global Virtual State
We compute the virtual global vector $\mathbf{g}$} by maintaining a buffer of the most recent previous affinity vectors, globally. Then, the virtual global vector is computed by performing aggregation over all the vectors in the buffer. The goal of the global vector is to detect a global trend (e.g, a new song or a new TV series that is globally streamed) before we are queried about a specific node. In practice, aggregating the buffer with the most recent vector selection is efficient and empirically effective, as we show later in Section˜4.
@@
Key Properties of NAViS. We note that NAViS generalizes EMA and other heuristics by allowing the gates ${\mathbf{z}}_{h}$ and ${\mathbf{z}}_{s}$ to adapt to new information, in contrast to a fixed $\alpha$. Large gate values enable the model to retain long-term information when beneficial. Notably, NAViS does not rely on neighbors’ hidden states, unlike other memory-based TGNNs, and therefore is compatible with the t-Batch mechanism (Kumar et al., 2019), enabling efficient batching without missing updates. In addition, we show in Appendix˜F that incorporating global information via $\mathbf{g}$} can improve the accuracy of the predicted affinity vector when global trends affect the nodes’ affinities.
NAViS for large-scale graphs. For graphs with $N$ nodes, the number of learnable parameters in NAViS scales as $O(N)$, which can be prohibitive for graphs with millions of nodes. To make NAViS practical at this scale, we introduce a sparsified affinity prediction pipeline. Specifically, for each node we retain only the entries corresponding to candidate target nodes. In real-world settings, such as streaming services where users and movies are nodes and interactions are edges, we are interested in each user’s affinity to movies rather than to other users. In practice, this substantially reduces the parameter count of NAViS. For example, on the tgbn-token dataset (Shamsi et al., 2022), which records user–token interactions, NAViS requires about (5,000) parameters, while the graph contains over (60,000) nodes. We further provide a detailed empirical runtime and memory analysis of NAViS in Appendix˜G.
@@ 3.3 Learning with Rank-Based Loss: Why Cross-Entropy Fails
With NAViS specified in Section˜3.2, we now turn to the question of how to train it effectively. Because most downstream uses of affinity vectors depend on the induced ordering of candidates rather than the actual affinity values (Huang et al., 2023; Tjandra et al., 2024), the choice of the loss function is critical. Most TGNNs use the cross-entropy loss (Luo and Li, 2022; Yu et al., 2023; Tjandra et al., 2024), which treats the output as a categorical distribution and ignores ordinal structure. As we show next, this property is suboptimal, and requires a designated loss to address this limitation.
@@
There exist infinitely many triplets of $\mathbf{y}$}, a ground-truth affinity vector, and ${\mathbf{s}}_{1}$, ${\mathbf{s}}_{2}$, two predicted logits of affinity vectors such that: ${\mathbf{s}}_{1}$ ordering exactly matches the ground-truth node affinity scores (i.e., yields a perfect ordering), ${\mathbf{s}}_{2}$ differs from ${\mathbf{s}}_{1}$ only in the logit of the most-affinitive node or the least-affinitive node—thereby making the ordering of ${\mathbf{s}}_{2}$ different from ${\mathbf{s}}_{1}$ and $\mathbf{y}$}, and $\ell_{\mathrm{CE}}({\mathbf{s}}_{1},{\mathbf{y}})>\ell_{\mathrm{CE}}({\mathbf{s}}_{2},{\mathbf{y}})$.
@@
Set ${\mathbf{y}}=[0.4,0.6]$, ${\mathbf{s}}_{1}=[1,2.4]$ (correctly rank logits), ${\mathbf{s}}_{2}=[1,0.6]$ (wrongly ranked logits). Then $0.75=\ell_{\mathrm{CE}}({\mathbf{s}}_{2},{\mathbf{y}})<\ell_{\mathrm{CE}}({\mathbf{s}}_{1},{\mathbf{y}})=0.78$, where ${\mathbf{s}}_{1}$ ranks the same as $\mathbf{y}$} while ${\mathbf{s}}_{2}$ does not. Since $\ell_{\mathrm{CE}}$ is a continuous function, there are infinitely many such triplets. ∎
@@
Pairwise Margin Regularization. In our experiments, we discovered that the use of the loss in Equation˜16 alone is not sufficient for future node affinity prediction tasks, as we elaborate in Appendix˜F. Hence, we suggest the following regularization:
@@
Here, $\Delta$ is a hyperparameter that represents the minimum margin required between each pair of affinity scores. The goal of this regularization is to prevent NAViS from incorrectly learning to shrink the affinity scores to minimize Equation˜16. In Appendix˜H we show that the undesirable property in Theorem˜3 does not hold for the suggested loss function.
@@ 4 Experiments
We evaluate NAViS across multiple future node affinity prediction benchmarks and compare it with recent state-of-the-art baselines, including both heuristics and TGNNs. Sections˜4.1 and 4.2 present our key empirical findings, and Appendix˜F provides additional ablation studies. Our experiments aim to address the following research questions: (RQ1) How does NAViS perform compared to prior art for future node affinity prediction? (RQ2) Does our method generalize across various types of graphs? (RQ3) What is the contribution of each component in NAViS?
Experimental setup We compare NAViS to the following TGNN baselines JODIE(Kumar et al., 2019), TGAT(Xu et al., 2020), CAWN (Wang et al., 2021b), TCL (Wang et al., 2021a), GraphMixer(Cong et al., 2023),DyGFormer(Yu et al., 2023),DyRep(Trivedi et al., 2019), TGN(Rossi et al., 2020), TGNv2(Tjandra et al., 2024) and the standard heuristics presented in (Huang et al., 2023) and in (Tjandra et al., 2024). Following the standard protocols (Huang et al., 2023), we use a 70%-15%-15% chronological split, train for 50 epochs, use a batch size of 200, and report the average NDCG@10 (Järvelin and Kekäläinen, 2002) over three runs. We include both future node affinity prediction settings: (1) using the full fine-grained CTDG up to the prediction time, and (2) using only previous ground-truth affinity vectors. Previous TGNNs only support the first setting (Tjandra et al., 2024), and, therefore, only heuristics are included in the second setting comparisons.